Show that the function f:]-1,1] [ --> R given by f(x) =x^3 is uniformly continuous and deduce that f is continuous at the point zero?
1
Expert's answer
2019-09-30T11:00:08-0400

Remember that

y= f(x) is uniformly continuous  on ]-1,1] if "\forall \epsilon >0 \forall x,y \in ]-1,1] \vert x-y\vert< \delta \Rightarrow \vert f(x)-f(y)\vert< \epsilon" .


If "\forall x,y \in ]-1,1]" :

"|x^3 - y^3| = |(x - y) ( x^2 + yx + y^2)|= \\=|x - y| | x^2 + yx + y^2| \leq\\\leq |x - y|( x^2 + | yx| + y^2) <\delta |1+1+1| =3\delta"

Proof.

Let "\epsilon >0 ." Choose "\delta= \frac{\epsilon}{3}"  .

Then "\forall x,y \in ]-1,1]" with "\vert x-y\vert< \delta" we have

"|x^3 - y^3| = |(x - y) ( x^2 + yx + y^2)|= \\=|x - y| | x^2 + yx + y^2| \leq \\ \leq |x - y|( x^2 + | yx| + y^2) <\delta |1+1+1| =3\delta=3\frac{\epsilon}{3}=\epsilon\\" .


If "\vert x-y\vert< \delta" :"|x^3 - y^3| <\epsilon" .

So

"f(x)=x^3" is uniformy continuous on ]-1,1] .


Since f(x) is uniformy continuous ]-1,1]  then it  is continuous at the point zero (x=0).